OpenStudy (anonymous):

If a gas-filled balloon rises 40 feet in the 1st minute, 60 feet in the 2nd minute, and 80 feet in the 3rd minute, how many feet will the balloon rise during the 15th minute? A. 300 feet B. 320 feet C. 340 feet D. 360 feet E. 380 fee

OpenStudy (lucaz):

you have a pattern 40 to 60 = 20, 60 to 80 = 20 1st min = 60 2nd min = 60 + 20 15th min = 60 +(14 x 20)

OpenStudy (anonymous):

b